# TFusionOcc: Student’s t-Distribution Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### v1

@@ III-C Camera-based Pseudo 3D Point Cloud and Multi-Scale Cam Dense Depth Maps Generation
The pseudo 3D point cloud $Pt^{cam}$ and camera-based multi-scale dense depth maps $V^{cam}_{depth}$ are generated through the DepthNet module as shown in Figure 3. The DepthNet module takes $V^{Cam}_{Encode}$ as input and leverages a SECONDFPN [32] to generate a single-scale visual feature $V^{cam}_{sec}=\left\{V_{n}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, whose resolution is $\frac{1}{8}$ of the input image. The $V^{cam}_{sec}$ is further passed to an MLP to estimate the ray depth per-pixel, resulting in a single-scale depth map $V_{depth\_n}\in\mathbb{R}^{N\times D\times H_{l}\times W_{l}}$. Bilinear interpolation with a 0.5 downsampling factor is applied iteratively to $V_{depth\_n}$, resulting in two additional depth maps that have $\frac{1}{16}$ and $\frac{1}{32}$ resolution of the input image. Instead of performing depth estimation on each scale visual feature, our approach saves a certain amount of computation overhead while still maintaining a relatively high depth estimation precision. The predicted depth map and the two extra downsampled depth maps form $V_{depth}^{cam}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Meanwhile, the pseudo 3D point cloud $Pt^{cam}$ is acquired only from $\frac{1}{8}$ scale depth map $V_{depth\_n}$. Each pixel ray has a set of depth bins separated by 0.5m intervals. We followed the approach proposed in [7] to use 3D semantic occupancy annotations to supervise depth estimation. For each pixel ray, we take the highest probability depth bins of $K$ to form $Pt^{cam}$.
@@ III-D Skeleton Merge Module
The primitive anchor initialization process is shown in Figure 4. Given the pseudo 3D point cloud derived from the camera $Pt^{cam}$ and the real 3D point cloud of the lidar $Pt^{lidar}$, we first apply the cylindrical partition to both point clouds, resulting in two volumes of voxels $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ under cylindrical coordinates. We then predefine the total primitive anchors of $M+N$, where the anchors $M$ are sampled from $Vol^{Lidar}_{Cylin}$ and the anchors $N$ are sampled from $Vol^{Cam}_{Cylin}$, with a ratio of $M:N=3:1$. Following [37, 7], we adopt the farthest point sampling strategy to obtain $M$ Lidar anchors from $Vol^{Lidar}_{Cylin}$, which act as the main structural skeleton. Using these Lidar anchors as reference, any voxel in $Vol^{Cam}_{Cylin}$ that spatially overlaps with the LiDAR anchors is removed. In addition, we impose a radius constraint r=5r=5m to perform range filtering. Any voxel in $Vol^{Cam}_{Cylin}$ whose closest Lidar anchor is farther away than $r$ is also discarded. Then, $N$ Camera anchors are obtained using the farthest point sampling strategy on the remaining voxels in $Vol^{Cam}_{Cylin}$. Based on the aforementioned filtering and sampling strategy, we acquire a final set of $M+N$ anchors whose locations will be used to initialize a set of T-primitives.
The skeleton merge module takes advantage of the lidar strength in preserving the fine-grained 3D geometry of the scene and takes advantage of its $Vol^{Lidar}_{Cylin}$ to extract the main skeleton. Meanwhile, the module also leverages the camera’s strength in handling occlusion scenes and uses its $Vol^{Cam}_{Cylin}$ to extract the augmented skeleton surrounding the main skeleton. The fused skeleton preserves the finer-grained 3D geometry of the scene by combining the advantages of both modalities.
III-E Multi-scale-Fused Dense Depth-Maps Generation
Figure 5: Multi-Scale Fused Dense Depth Maps Generation. The upper part exhibits general multi-modality depth map fusion from a depth map perspective, and the bottom part demonstrates the detailed multi-modality depth map fusion from a per-pixel ray perspective.
Given multi-scale camera dense depth maps $V^{cam}_{depth}$ and multi-scale lidar sparse depth maps $V_{depth}^{lidar}$, we first go through per-pixel level value addition followed by a value clamp, resulting in the final multi-scale fused dense depth maps as shown in Figure 5 upper. The detail of the addition of the per-pixel ray value is shown in Figure 5 bottom. For the camera side, each pixel ray has $D$ number of depth bins, and the interval between each depth bin is 0.5m. The DepthNet predicts the probability that each depth bin being occupied with a value in the range of [0, 1]. On the lidar side, each pixel ray also has $D$ depth bins, but only one depth bin has been assigned a probability value of 1.0, while all other depth bins are assigned a probability value of 0. When adding the camera pixel ray to the lidar pixel ray, the resulting fused pixel ray may have some depth-bin probability values exceeding 1.0; therefore, these values need to be clamped within the range [0, 1].
This multi-modality depth map fusion, on the one hand, solves the sparsity problem of lidar-based depth maps, and on the other hand, mitigates the inaccurate depth estimation problem of camera-based depth maps, resulting in denser and highly precise fused depth maps.
III-F Early-Fusion Module
@@
Figure 6: Early-Fusion Module. Each occupied voxel centre in $Vol^{Lidar}_{Cylin}$ serves as an anchor and is projected onto multi-scale visual features to aggregate semantic information, yielding a semantic-aware lidar feature.
@@
The detailed inner structure of the Early-Fusion module is shown in Figure 6. Given the sparse volume of the voxel $Vol_{Cylin}^{Lidar}$ in cylindrical coordinates, we leverage each occupied voxel center as an anchor and convert each anchor position from cylindrical coordinates to Cartesian coordinates. Then, each anchor is projected onto multi-scale visual features $V^{Cam}$ to aggregate semantic information at three scales. The aggregated three-scale visual features are weighted, summed, and concatenated back to the corresponding voxel feature, then fed through an MLP layer to perform the feature fusion, resulting in the last semantic-aware lidar feature $Vol_{sem}^{Lidar}$.
III-G T-Primitive Representation
@@
Following the GaussianFormer [9, 7, 40] series, we propose the probabilistic Student’s T-distribution superposition as an efficient and effective 3D scene representation. Specifically, we propose three types of primitives that are enhanced by the inner kernel of T-distribution or T-distribution-like. The first is the Student-T-distribution-based primitive, which has the property as follows:
@@
where $m_{i}$, $s_{i}$, $r_{i}$, $\alpha_{i}$ and $c_{i}$ refer to the mean, scale value for the $x,y,z$ axes, quaternion vector, opacity and semantic vector of the $i$-th primitive, respectively. The second is the Student-T-distribution-based superquadric primitive, which has the property as follows:
$P_{SQi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1\_i},\epsilon_{2\_i},c_{i})$ (4)
@@
where $\epsilon_{1}$ and $\epsilon_{2}$ are the shape exponents of the superquadric. The third is the Student-T-distribution-based superquadric with inverse-warp primitive that has the property as follows:
$P_{SQIWi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1\_i},\epsilon_{2\_i},w_{i},c_{i})$ (5)
where $\omega_{i}$ is the learnable weighting scaler valued between -1 and 1.
Similarly to [7], we decompose the 3D modeling target into geometry and semantics predictions, and adopt the multiplication theorem of probability and the T-mixture model to address them probabilistically, respectively. What is more, compared to the Gaussian mixture model, the T-mixture model significantly improves robustness against outliers.
III-G1 Student-T-Distribution Based Geometry Prediction
@@
To restrict T-Primitives to represent only occupied regions for geometry prediction, we interpret the T-distribution primitives $P^{T}=\left\{p^{T}_{i}\right\}_{i=1}^{K}$ as the probability of their surrounding space being occupied. In detail, we assign a probability value of 100% to the centers of the T-distribution, which decays with respect to the distance from the centers $m$:
$\alpha\left(x;p^{T}_{i}\right)=\left[1+\frac{1}{\nu}\left(x-m\right)^{T}\Sigma^{-1}\left(x-m\right)\right]^{-\frac{\nu+3}{2}}$ (6)
where $\nu$ refers to the degree of freedom, it is obtained by calculating how many voxels are intercepted by each T-primitive minus 1, and $\alpha\left(x;p^{T}_{i}\right)$ denotes the probability of the point x being occupied induced by the $i$-th T-primitive. Equation 6 assigns a high probability of occupancy when the point x is close to the center $m$ of the $i$-th T-primitive, which prevents any T-primitive from describing an empty area. To further derive the overall probability of occupancy, we assume that the probabilities of a point being occupied by different T-primitives are mutually independent, and thus we can aggregate them according to the multiplication theorem of probability:
$\alpha\left(x\right)=1-\prod_{i=1}^{K}\left(1-\alpha\left(x;p^{T}_{i}\right)\right)$ (7)
where $\alpha(x)$ represents the overall probability of occupancy at point x.
III-G2 T-Mixture Model Based Semantic Prediction
In addition to the Student-T-distribution-based Geometry modeling, we also need to perform the Student-T-distribution-based semantics prediction. Following [7], we interpret the set of T-primitives as a T-mixture model, where the semantics prediction could be formulated as calculating the expectation of semantics given the probabilistic T-mixture model.
To be specific, we take the original opacity properties $\alpha$ as the prior distribution of the T-primitives, which is normalized by $l1$. Then, the T-primitive probabilistic distribution is adopted as a conditional probability. In addition, we normalize each T-primitive’s original semantics properties $c_{i}$ with a softmax activation function, yielding $\tilde{c}_{i}$ to ensure the boundness of predicted semantics. Lastly, the expectation calculation of the T-mixture model can be formulated as:
$e\left(x;P^{T}\right)=\sum_{i=1}^{K}p\left(P^{T}_{i}|x\right)\tilde{c}_{i}=\frac{{\textstyle\sum_{i=1}^{K}p\left(x|P^{T}_{i}\right)a_{i}\tilde{c}_{i}}}{{\textstyle\sum_{j=1}^{K}p\left(x|P^{T}_{j}\right)a_{j}}}$ (8)
where $p\left(P^{T}_{i}|x\right)$, $p\left(x|P^{T}_{i}\right)$ and $\tilde{c}_{i}$ denote the posterior probability of point x belonging to the $i$-th T-primitive distribution, the conditional probability of point x given the $i$-th T-primitive distribution, and the softmax-normalized semantic properties, respectively.
Lastly, based on the geometry and semantic predictions, we combine the two prediction results together, resulting in the final 3D semantic occupancy prediction. The combination process is formulated as follows.
$Occ=\left[1-\alpha\left(x\right);\alpha\left(x\right)\cdot e\left(x;P^{T}\right)\right]$ (9)
where $1-\alpha\left(x\right)$ refers to the probability of the empty class and $\alpha\left(x\right)\cdot e\left(x;P^{T}\right)$ refers to the final semantic probabilities. The $\alpha\left(x\right)$ is used to weight the initial semantic prediction.
III-G3 Student-T-Distribution Based Primitive
@@
The straightforward T-distribution primitive $p\left(x|P^{T}_{i}\right)$ can be formulated as:
$p\left(x|P^{T}_{i}\right)$ $=\frac{\Gamma\left(\frac{\nu+3}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\cdot\left(\nu\pi\right)^{3/2}\cdot\left|\Sigma\right|^{1/2}}\cdot$
$\left[1+\frac{1}{\nu}\left(x-m\right)^{T}\Sigma^{-1}\left(x-m\right)\right]^{-\frac{\nu+3}{2}}$ (10)
where $\nu$ refers to the degree of freedom, the same as in Equation 6, it is obtained by calculating how many voxels are intercepted by each T-primitive minus 1. $\Gamma$ refers to the gamma function, and performs the $\Gamma\left(n\right)=\left(n-1\right)!$ operation. Compared to Gaussian-primitive, this T-primitive is highly tailed in its probability distribution, leading to better robustness against outliers.
III-G4 Student-T-Distribution Based SuperQuadric Primitive
@@
Following the quadricformer [40], we use the T-probabilistic modeling mechanism to convert superquadrics into occupancy probabilities. We first transform the 3D point position $(X,Y,Z)$, which is occupied by the T-Superquadric primitive, into its local coordinate system as:
$\begin{bmatrix}X_{S}\\ Y_{S}\\ Z_{S}\\ \end{bmatrix}=R\times\begin{bmatrix}X-m_{x}\\ Y-m_{y}\\ Z-m_{z}\\ \end{bmatrix}$ (11)
@@
where $(X_{S},Y_{S},Z_{S})$ refers to the local coordinate under the T-Superquadric primitive, and R is the rotation matrix of each T-Superquadric primitive. Then, the occupancy probability of the 3D point $(X,Y,Z)$ associated with each T-Superquadric primitive can be computed as:
$f\left(X_{S},Y_{S},Z_{S}\right)=\left(\left(\frac{X_{S}}{S_{x}}\right)^{\frac{2}{\epsilon_{2}}}+\left(\frac{Y_{S}}{S_{y}}\right)^{\frac{2}{\epsilon_{2}}}\right)^{\frac{\epsilon_{2}}{\epsilon_{1}}}+\left(\frac{Z_{S}}{S_{z}}\right)^{\frac{2}{\epsilon_{1}}}$ (12)
$p\left(x|P^{T}_{SQi}\right)=\left(1+\frac{1}{\nu}\cdot f\left(X_{S},Y_{S},Z_{S}\right)\right)^{-\frac{\nu+3}{2}}$ (13)
@@
where $(S_{X},S_{Y},S_{Z})$ refers to the scale parameters along the X, Y, Z axes, $\epsilon_{1}$ and $\epsilon_{2}$ are the shape exponents of the superquadric.
@@
III-G5 Student-T-Distribution Based SuperQuadric With Inverse-Warp Primitive

### v2

@@ III-B Camera-based Pseudo 3D Point Cloud and Multi-Scale Cam Dense Depth Maps Generation
The pseudo 3D point cloud $Pt^{cam}$ and camera-based multi-scale dense depth maps $V^{cam}_{depth}$ are generated through the DepthNet module as shown in Fig. 3. The DepthNet module takes $V^{Cam}_{Encode}$ as input and leverages a SECONDFPN [33] to generate a single-scale visual feature $V^{cam}_{sec}=\left\{V_{n}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, whose resolution is $\frac{1}{8}$ of the input image. The $V^{cam}_{sec}$ is further passed to an MLP to perform the multi-hot per-ray depth estimation, resulting in a depth map $V_{depth\_1}\in\mathbb{R}^{N\times D\times H_{l}\times W_{l}}$ with each pixel ray having a set of depth bins separated by 0.5m intervals, and each depth bin on the pixel ray has a probability range in $[0,1]$. Following [7], we adopt the 3D semantic occupancy label to supervise $V_{depth\_1}$. Then, bilinear interpolation with a 0.5 downsampling factor is applied iteratively to $V_{depth\_1}$, resulting in two additional depth maps that have $\frac{1}{16}$ and $\frac{1}{32}$ resolution of the input image. The predicted depth map and the two extra downsampled depth maps form $V_{depth}^{cam}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Meanwhile, the pseudo 3D point cloud $Pt^{cam}$ is acquired only from $\frac{1}{8}$ scale depth map $V_{depth\_1}$.
III-C Multi-scale-Fused Dense Depth-Maps Generation
Figure 4: Multi-Scale Fused Dense Depth Maps Generation. This demonstrates the detailed multi-source depth map fusion from a per-pixel ray perspective.
Given the multi-scale dense depth maps from cameras, $V^{cam}_{depth}$, and the multi-scale sparse depth maps from LiDAR, $V^{lidar}_{depth}$, we perform element-wise addition at the per-pixel level, followed by a value clamping operation. This process produces the final multi-scale fused dense depth maps, as illustrated in Fig. 4. All depth maps from different sensors share a consistent representation, with the same number of depth bins $D$ and a uniform interval of 0.5 m between adjacent bins along each pixel ray. However, their value distributions differ: on the camera side, $V^{cam}_{depth}$ is represented as multi-hot per-pixel ray depth maps, where each depth bin takes a value within the range $[0,1]$. In contrast, $V^{lidar}_{depth}$ is represented as single-hot per-pixel ray depth maps, where only one depth bin is assigned a value of 1.0. When the camera and LiDAR depth maps are combined via per-pixel addition, the resulting fused pixel rays may contain depth-bin values exceeding 1.0. To ensure valid probability values, all depth-bin responses are subsequently clamped to the range $[0,1]$.
This fusion strategy combines the complementary strengths of the two modalities. The dense camera depth maps provide rich structural coverage but are less accurate, whereas the LiDAR depth maps are sparse but geometrically precise. Their combination yields denser, more reliable fused depth maps, which are subsequently used to enhance visual features and facilitate cross-modal fusion.
@@ III-D Skeleton Merge Module
The primitive anchor initialization process is shown in Fig. 5. Given the pseudo 3D point cloud derived from the camera $Pt^{cam}$ and the real 3D point cloud of the LiDAR $Pt^{lidar}$, we first apply the cylindrical partition to both point clouds, resulting in two volumes of voxels $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ under cylindrical coordinates. We then predefine the total primitive anchors of $M+N$, where the anchors $M$ are sampled from $Vol^{Lidar}_{Cylin}$ and the anchors $N$ are sampled from $Vol^{Cam}_{Cylin}$, with a ratio of $M:N=3:1$. We adopt the farthest point sampling strategy to obtain $M$ LiDAR anchors from $Vol^{Lidar}_{Cylin}$, which act as the main structural skeleton. Using these LiDAR anchors as reference, any voxel in $Vol^{Cam}_{Cylin}$ that spatially overlaps with the LiDAR anchors is removed. In addition, we impose a radius constraint r=5r=5m to perform range filtering. Any voxel in $Vol^{Cam}_{Cylin}$ whose closest LiDAR anchor is farther away than $r$ is also discarded. Then, $N$ Camera anchors are obtained using the same sampling strategy on the remaining voxels in $Vol^{Cam}_{Cylin}$. Based on the aforementioned filtering and sampling strategy, we acquire a final set of $M+N$ anchors whose locations will be used to initialize a set of T-primitives.
This design exploits the complementary strengths of the two modalities: LiDAR provides the main geometric structure with higher spatial accuracy, while the camera branch supplements local details and occluded regions. The resulting fused skeleton, therefore, offers a better initialization for subsequent T-primitive modeling.
III-E Early-Fusion Module
@@
Figure 6: Early-Fusion Module. Each occupied voxel center in $Vol^{Lidar}_{Cylin}$ serves as an anchor and is projected onto multi-scale visual features to aggregate semantic information, yielding a semantic-aware LiDAR feature.
@@
The detailed inner structure of the Early-Fusion module is shown in Fig. 6. Given the sparse volume of the voxel $Vol_{Cylin}^{Lidar}$ in cylindrical coordinates, we leverage each occupied voxel center as an anchor and convert each anchor position from cylindrical coordinates to Cartesian coordinates. Then, each anchor is projected onto multi-scale visual features $V^{Cam}$ to aggregate semantic information at three scales. The aggregated three-scale visual features are weighted, summed, and concatenated back to the corresponding voxel-associated LiDAR feature, then fed through an MLP layer to perform the feature fusion, resulting in the last semantic-aware LiDAR feature $Vol_{sem}^{Lidar}$.
This module injects image semantics into sparse LiDAR features at an early stage, allowing the LiDAR branch to retain its geometric precision while gaining richer semantic context from the camera branch.
III-F T-Primitive Representation
@@
We propose a probabilistic superposition of Student’s t-distribution-based primitives as an efficient and effective 3D scene representation. Specifically, we propose three types of primitives that are enhanced by the inner kernel of the T-distribution or T-distribution-like. The first is the Student-T-distribution-based plain primitive, which has the attributes as follows:
@@
where $m_{i}$, $s_{i}$, $r_{i}$, $\alpha_{i}$ and $c_{i}$ refer to the mean, scale value for the $x,y,z$ axes, quaternion vector, opacity and semantic vector of the $i$-th primitive, respectively. The second is the Student-T-distribution-based superquadric primitive, which has the attributes as follows:
$P_{SQi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1_{i}},\epsilon_{2_{i}},c_{i})$ (4)
@@
where $\epsilon_{1}$ and $\epsilon_{2}$ are the shape exponents of the superquadric. The third is the Student-T-distribution-based superquadric with inverse-warp primitive that has the attributes as follows:
$P_{SQIWi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1_{i}},\epsilon_{2_{i}},\left\{\omega_{i,k}\right\}_{k=1}^{24},c_{i})$ (5)
where $\left\{\omega_{i,k}\right\}_{k=1}^{24}$ is a set of learnable weighting scalers used to determine the magnitude of applied deformation, and each $\omega_{i,k}$ has a value range between -1 and 1.
Inspired by [7], we perform 3D geometry prediction and semantic prediction separately and adopt the multiplication theorem of probability and the T-mixture model to address them probabilistically.
III-F1 Student-T-Distribution and Multiplication Theorem Based Geometry Prediction
@@
In practice, we interpret the T-distribution primitives $P^{T}=\left\{P^{T}_{i}\right\}_{i=1}^{K}$ as the probability of their surrounding space being occupied. In detail, we assign a probability value of 100% to the centers of the $P^{T}_{i}$, which decays with respect to the distance from the centers $m_{i}$. Assuming conditional independence of occupancy among different T-primitives, the final occupancy probability at x can be computed as:
$p_{o}\left(x\right)=1-\prod_{i=1}^{K}\left(1-p\left(x|P^{T}_{i}\right)\right)$ (6)
where $p\left(x|P^{T}_{i}\right)$ denotes the probability of point x being occupied induced by the $i$-th T-primitive. Depending on the primitive type, it is instantiated by (8), (12), or (16). $p_{o}\left(x\right)$ refers to the overall probability of occupancy at point x.
III-F2 T-Mixture Model Based Semantic Prediction
A set of T-primitives can be considered as a T-mixture model, and the final semantic predictions can be calculated via a weighted aggregation of semantic probabilities from all contributing T-primitives in the scene as:
$p_{sem}\left(x\right)=\frac{{\textstyle\sum_{i=1}^{K}p\left(x|P^{T}_{i}\right)a_{i}\tilde{c}_{i}}}{{\textstyle\sum_{j=1}^{K}p\left(x|P^{T}_{j}\right)a_{j}}}$ (7)
where $\tilde{c}_{i}$ represents the softmax-normalized semantic property of $i$-th T-primitive, and $a_{i}$ refers to $l1$ normalized opacity property of $i$-th T-primitive.
III-F3 Student-T-Distribution Based Plain Primitive
@@
The plain T-distribution primitive $p\left(x|P^{T}_{i}\right)$ with the lowest geometric flexibility can be formulated as:
$p\left(x|P^{T}_{i}\right)$ $=\frac{\Gamma\left(\frac{\nu_{i}+3}{2}\right)}{\Gamma\left(\frac{\nu_{i}}{2}\right)\cdot\left(\nu_{i}\pi\right)^{3/2}\cdot\left|\Sigma\right|^{1/2}}\cdot$
$\left[1+\frac{1}{\nu_{i}}\left(x-m\right)^{T}\Sigma^{-1}\left(x-m\right)\right]^{-\frac{\nu_{i}+3}{2}}$ (8)
where $\nu_{i}$ denotes the degrees of freedom of the $i$-th T-primitive. In our setting, we treat the number of voxels intersected by a T-primitive in the final 3D voxel grid as an approximation of its observable local support. Therefore, we adaptively define
$\nu_{i}=N_{i}-1,$ (9)
where $N_{i}$ is the number of intersected voxels associated with the $i$-th T-primitive. In this way, primitives supported by fewer voxels retain heavier tails for better robustness, while primitives supported by more voxels yield more concentrated distributions. This adaptive design allows each T-primitive to adjust its probabilistic shape according to its local geometric support in the scene. $\Gamma\left(\cdot\right)$ denotes the gamma function, which satisfies $\Gamma\left(n\right)=\left(n-1\right)!$ for positive integers $n$. Compared to the plain Gaussian-primitive, this T-primitive is highly tailed in its probability distribution, leading to better robustness against outliers.
III-F4 Student-T-Distribution Based SuperQuadric Primitive
@@
The T-probabilistic modeling mechanism is adopted to convert superquadrics into occupancy probabilities. We first transform the 3D point position $(x,y,z)$, which is occupied by the $i$-th T-Superquadric primitive, into its local coordinate system as:
$\begin{bmatrix}X_{S_{i}}\\ Y_{S_{i}}\\ Z_{S_{i}}\\ \end{bmatrix}=R_{i}\times\begin{bmatrix}x-m_{x_{i}}\\ y-m_{y_{i}}\\ z-m_{z_{i}}\\ \end{bmatrix}$ (10)
@@
where $(X_{S_{i}},Y_{S_{i}},Z_{S_{i}})$ refers to the local coordinate under the $i$-th T-Superquadric primitive, and $R_{i}$ is the rotation matrix of the $i$-th T-Superquadric primitive. Then, the occupancy probability of the 3D point $(x,y,z)$ associated with the $i$-th T-Superquadric primitive can be computed as:
$f\left(X_{S_{i}},Y_{S_{i}},Z_{S_{i}}\right)=\left(\left(\frac{X_{S_{i}}}{S_{x_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}+\left(\frac{Y_{S_{i}}}{S_{y_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}\right)^{\frac{\epsilon_{2_{i}}}{\epsilon_{1_{i}}}}+\left(\frac{Z_{S_{i}}}{S_{z_{i}}}\right)^{\frac{2}{\epsilon_{1_{i}}}}$ (11)
$p\left(x|P^{T}_{SQi}\right)=\left(1+\frac{1}{\nu_{i}}\cdot f\left(X_{S_{i}},Y_{S_{i}},Z_{S_{i}}\right)\right)^{-\frac{\nu_{i}+3}{2}}$ (12)
@@
where $(S_{x_{i}},S_{y_{i}},S_{z_{i}})$ refers to the scale parameters of $i$-th T-Superquadric primitive along the $X,Y,Z$ axes and $\nu_{i}$ refers to the degree of freedom, and has the same definition as in (8).
@@
III-F5 Student-T-Distribution Based SuperQuadric With Inverse-Warp Primitive